# Understanding and Improving Length Generalization in Hierarchical Sparse Attention Models

### v1

@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) has catalyzed an urgent demand for models capable of processing long contexts (Brown et al., 2020; Touvron et al., 2023; OpenAI et al., 2024). However, the performance of standard Transformer architectures (Vaswani et al., 2023), which serve as the universal backbone, is well-documented to degrade sharply on out-of-distribution context lengths (Hsieh et al., 2024; Xiao et al., 2024; Lu et al., 2024; Kuratov et al., 2024). Concurrently, naively increasing the context window during training is often infeasible due to the quadratic complexity of the attention mechanism, which leads to prohibitive computational costs. This has spurred the community to investigate novel architectures designed explicitly for length extrapolation.
We posit that ideal length extrapolation necessitates two key properties: (1) the ability to train on shorter sequences while achieving stable or even lower perplexity on longer sequences during inference, and (2) the capacity to effectively utilize the entire context. Significant research has focused on the first property. For instance, architectures like Sliding Window Attention (Beltagy et al., 2020) and recurrent models can maintain a nearly constant perplexity as the context extends beyond the training length (Ruiz & Gu, 2025). However, these methods fall short on the second property. The former is constrained by a fixed-size local receptive field, while the latter compresses the entire history into a fixed-size state, creating an information bottleneck. As revealed by in-context retrieval tasks like RULER(Hsieh et al., 2024), the inability to access and utilize arbitrary information from long contexts remains a critical limitation for many applications, from multi-round chatbots to agent-based systems (Liu et al., 2023; Pan et al., 2025; Guo et al., 2023).
@@ 2 Related Work
Approaches to long-context modeling can be broadly categorized. The first comprises local methods like Sliding Window Attention (Beltagy et al., 2020; Xiao et al., 2024). While effective at maintaining perplexity, they are fundamentally limited by a fixed-size local receptive field, preventing access to information beyond their local window. A second category includes recurrent architectures  (Gu & Dao, 2024; Dao & Gu, 2024; Sun et al., 2023; Qin et al., 2023; Peng et al., 2023a; Yang et al., 2025). Recent work shows that zero-initialized recurrent models struggle to generalize to longer sequences; it attributes the failure to models entering unexplored states and introduces State Passing and Truncated Backpropagation Through Time (Sutskever, 2013; Williams & Peng, 1990) to improve length generalization, achieving stable perplexity on out-of-distribution lengths (Ruiz & Gu, 2025). However, on challenging retrieval tasks (Hsieh et al., 2024), performance still drops significantly beyond the training length. This is because the entire context history is compressed into a fixed-size state, sacrificing the ability to recall specific, distant information losslessly. The third category seeks to extend full attention by modifying positional encodings (Press et al., 2022; Peng et al., 2023b; Chen et al., 2023; Tan et al., 2025). Their training-free extrapolation gains are modest, typically limited to a small multiple of training length.
In pursuit of more scalable extrapolation, chunk-based sparse attention has emerged as a distinct and highly effective paradigm. Landmark Attention (Mohtashami & Jaggi, 2023) pioneered this by using dedicated landmark tokens to represent chunks for retrieval. While innovative, its length exptrapolation capability is limited to about 64x training length. NSA (Yuan et al., 2025) proposed an efficient hardware-aligned implementation of chunk selection but has very limited length extrapolation (Hu et al., 2025a). State-of-the-art training-free extrapolation has been achieved by Differentiable Retrieval-based Transformers (DRT) (Hu et al., 2025b) and RAMba (Hu et al., 2025a), which can generalize up to 1000x their training length on simple retrieval tasks by introducing a learnable hierarchical chunkwise sparse attention for retrieval. These methods also differ in their retrieval frequency, with DRT operating per-chunk while RAMba, NSA, and Landmark retrieve per-token. Our work provides a systematic analysis to dissect and identify the specific architectural components responsible for the extreme generalization capabilities of this emerging paradigm.
@@ 3.1 Random Context Access: The Key Ability to Utilize Full Context
Under this definition, different architectures exist on a spectrum of capability. The standard Transformer, with its full-attention mechanism, perfectly exemplifies RCA by granting unrestricted access to the entire KV cache. In stark contrast, many efficient architectures trade this property for computational gains by imposing structural constraints. These include: locality constraints, in which models like Sliding Window Attention are confined to a local neighborhood; information bottlenecks, created by State Space Models (SSMs) that compress history into a lossy, fixed-size state; and static access patterns (e.g., Child et al. (2019); Beltagy et al. (2020)), where data-independent sparse attention predetermines information pathways, with more details in Apx. A.1. These architectural bottlenecks fundamentally limit a model’s capacity for true long-range reasoning.
@@ 3.3 Hierarchical Sparse Attention and Model Design
Figure 1: DRT Architecture. The DRT (Hu et al., 2025b) model is primarily composed of Transformers with Sliding Window Attention (SWA), with layers divided into upper layers and lower layers. Lower decoders are composed of SWA followed by an FFN. At the architectural midpoint, a chunking layer splits representations from lower decoders into chunks, and encode them by chunk to build a global memory, each chunk of which contains a landmark vectors ($\mathbf{lmk}$) and encoded chunks ($\mathbf{E}$). Upper decoders each contains a local self-attention layer for local context, followed by an HSA to attend selected global memory and then an FFN.
As shown in Figure 1, the DRT model treats the global memory in a chunkwise approach. During decoding, the chunking layer gathers hidden representations into a buffer and encodes them as complete chunks. To maintain causality, the chunk containing the current token, as well as all subsequent chunks, are not accessible for retrieval. Following the notation of Sun et al. (2023), we use indices with $\cdot$] to indicate that it is indexed by chunk rather than by token. For instance, $\mathbf{H}_{[t]}:=\mathbf{h}_{t\cdot C+1:(t+1)\cdot C}\in\mathbb{R}^{C\times d}$ is the chunked hidden states. At chunk selection layer, a query vector $\mathbf{q_{t}}\in\mathbb{R}^{d_{\text{Rtrv}}}$ computes a dot-product similarity score on landmark $\mathbf{lmk}_{[i]}\in\mathbb{R}^{d_{\text{Rtrv}}}$, formulated as $s_{t,i}=\mathbf{q_{t}}\cdot\mathbf{{lmk}}_{[i]}\in\mathbb{R}$. The indices of the top-$N$ chunks are then selected, forming set $\mathcal{I}_{t}$.
A critical design choice lies in how these selected chunks are weighted, as this determines how the retrieval signal is backpropagated. The weighting function $w_{t,i}$ is zero for all non-selected chunks. For the selected chunks, the weights are assigned according to distinct strategies for NSA (Eq. 1a), DRT (Eq. 1b), and RAMba (Eq. 1c). Let $\mathcal{S}_{t}=\{s_{t,i}\}_{i\in\mathcal{I}_{t}}$.
This comparison highlights a clear progression: from the implicit binary weighting of NSA, to the proportional softmax weighting in DRT, and finally to the recency-biased allocation, stick-breaking process  [^1] Specifically, let $\mathcal{I}^{\prime}_{t}=\text{sort}(\mathcal{I}_{t})$ be the selected chunk indices sorted by their scores, then $w_{t,k}=\sigma(s_{t,\mathcal{I}^{\prime}_{t,k}})\prod_{j<k}(1-\sigma(s_{t,\mathcal{I}^{\prime}_{t,j}}))$ where $\sigma$ is the sigmoid function.  (Tan et al., 2025) used in RAMba.
@@ 4.1 Chunk Information preprocess
Chunked Hidden States: Disentangling Representations for Retrieval and Prediction. Considering the dual purposes of $\mathbf{h}^{\frac{L}{2}}_{t}$, we posit that simply operating on these raw hidden states conflates these two distinct functional objectives. The introduction of non-linear layers aims to disentangle these two roles. By processing the raw states, these components can produce refined representations specifically optimized for retrieval, separate from the immediate demands of next-token prediction. A RMSNorm provides a simple mechanism to introduce non-linearity, while the Encoder takes this a step further by explicitly learning to re-encode the chunk’s information.
@@ 5 Experiments
5.1 Experimental Setup
Figure 3: BabiLong Evaluation Results: Our best DRT model demonstrates robust long-context extrapolation, maintaining high accuracy up to 8M tokens. In stark contrast, baseline models exhibit clear limitations. Full-attention methods collapse to near-zero accuracy shortly after their training length. Prior sparse methods like Landmark Attention show better generalization but also fail decisively at longer contexts ($\geq$64k). Sliding Window Attention plateaus at a low accuracy consistent with random guessing, while Mamba2’s performance degrades substantially.
Our experiments are conducted on a series of models based on the DRT architecture detailed in Figure 1. All model variants contain approximately 240M parameters and are trained with a context length of 4K. For our analysis, we replace the original GCA mechanism with our HSA module, utilizing a stick-breaking weighting strategy (Eq. 1c) to enable robust per-token retrieval.
Key hyperparameters for this setup are fixed across experiments. The sliding window size for the lower layers is 512 tokens, and the HSA chunk size is 64. Unless otherwise noted, we use a Top-K of 8 for chunk selection, giving the HSA module an effective receptive field of 512 tokens. A comprehensive description of our training recipes and further hyperparameter details is provided in Appendix B.
To complement this and more directly assess a model’s core retrieval ability in isolation, we also use RULER. From RULER, we select three representative sub-tasks: Single-Needle (S-N), Multi-Query Needle (MQ-N), and Variable Tracking (VT). Collectively, these tasks provide a targeted evaluation of a model’s capacity for precise, long-range information retrieval and tracking simple logical chains. As RULER requires accurately retrieve several tokens, it provides a strict test of models’ random context access ability.
Baselines. We compare our method against several baselines. Our primary full-attention baseline is Llama with YaRN and Mamba2 with State Passing (Ruiz & Gu, 2025), which is trained on a 4k context length. To ensure a fair comparison with other efficient attention mechanisms, we include Landmark Attention, configured to match our model’s training FLOPs during training by using a sequence length of 756. Furthermore, to benchmark against methods with a similar computational budget, we include Stick-breaking Attention and Sliding Window Attention. Both of these are trained with a 1k context window to keep the receptive field same to our approach.
5.2 Length Extrapolation Results
Table 2: RULER evaluation results. Our DRT best model maintains high accuracy up to 32M tokens, significantly outperforming all baselines in long-context scenarios. In Model Specs, ‘Enc’ denotes the number of encoder layers, ‘CLS’ indicates the use of a CLS token, and ‘Bypass’ refers to the Bypassing Residual Path.
@@ 5.2 Length Extrapolation Results
DRT 0 no no 72.28 21.78 3.96 32.67 61.39 6.93 1.98 23.43 61.80 1.12 0.00 20.97
DRT 1 no no 72.73 58.18 72.73 67.88 69.09 34.55 58.18 53.94 70.91 47.27 43.64 53.94
DRT 1 yes no 63.64 40.00 54.55 52.73 69.09 38.18 49.09 52.12 54.55 21.82 30.91 35.76
DRT 2 no no 63.37 63.37 72.28 66.34 58.42 48.51 66.34 57.76 67.42 42.70 70.79 60.30
DRT 2 yes no 83.17 46.53 82.18 70.63 78.22 28.71 80.20 62.38 82.02 24.72 61.80 56.18
DRT 0 no yes 80.20 64.36 45.54 63.37 71.29 46.53 15.84 44.55 82.02 31.46 3.37 38.95
DRT 1 no yes 69.09 56.36 45.45 56.97 58.18 30.91 10.91 33.33 63.64 36.36 12.73 37.58
DRT 1 yes yes 89.09 69.09 80.00 79.39 72.73 78.18 76.36 75.76 85.45 69.09 69.09 74.54
DRT 2 no yes 85.15 74.26 74.26 77.89 78.22 65.35 65.35 69.64 82.02 66.29 41.57 63.29
DRT 2 yes yes 88.12 76.24 93.07 85.81 79.21 77.23 90.10 82.18 82.02 68.54 88.76 79.77
@@
As shown in Figure 3, the results on BabiLong highlight the stark differences in extrapolation capability across architectures. The benchmark’s classification-like nature establishes a random-guess baseline around 0.3 accuracy, a level to which Sliding Window Attention predictably converges due to its limited context. Other baselines also fail to generalize, with full-attention and Landmark Attention collapsing at various points beyond their training regimes. Crucially, the figure also reveals the importance of our key design choices within the DRT framework. The DRT variants with a learnable chunk encoder consistently outperform those without.
@@
To further validate these findings and provide a more granular analysis of retrieval performance, we turn to the RULER benchmark, with detailed results in Table 2. The trends observed here strongly corroborate our conclusions from BabiLong. First, comparing our DRT models against the baselines reveals a significant performance gap in extrapolation. Llama with Yarn (Llama-Yarn), despite its strong in-domain performance at 4K, fails catastrophically at 32K, with their scores dropping to zero. Stick-breaking Attention (Stick-Break) and Landmark Attention demonstrate some extrapolation ability but exhibit significant performance decay and are consistently outperformed by our DRT models on longer contexts.
@@
Second, while all DRT variants showcase robust long-context performance, the results also highlight a clear performance hierarchy among our different architectural configurations. Our full model (DRT with Enc=2, CLS=yes, Bypass=yes) consistently achieves state-of-the-art results. It not only secures the highest average score of 87.92% at the 4K training length but also degrades remarkably gracefully, maintaining an average score of 79.77% even at an extreme context length of 32M. This indicates that the specific combination of our proposed components is critical for unlocking maximum performance, an interaction we will analyze in detail in the subsequent section.

### v2

@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) has catalyzed an urgent demand for models capable of processing long contexts (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2024). However, the performance of standard Transformer architectures (Vaswani et al., 2023), which serve as the universal backbone, is well-documented to degrade sharply on out-of-distribution context lengths (Hsieh et al., 2024; Xiao et al., 2024; Lu et al., 2024; Kuratov et al., 2024). Concurrently, naively increasing the context window during training is often infeasible due to the quadratic complexity of the attention mechanism, which leads to prohibitive computational costs. This has spurred the community to investigate novel architectures designed explicitly for length extrapolation.
We posit that ideal length extrapolation necessitates two key properties: (1) the ability to train on shorter sequences while achieving stable or even lower perplexity on longer sequences during inference, and (2) the capacity to effectively utilize the entire context. Significant research has focused on the first property. For instance, architectures like Sliding Window Attention (Beltagy et al., 2020) and recurrent models can maintain a nearly constant perplexity as the context extends beyond the training length (Buitrago & Gu, 2025). However, these methods fall short on the second property. The former is constrained by a fixed-size local receptive field, while the latter compresses the entire history into a fixed-size state, creating an information bottleneck. As revealed by in-context retrieval tasks like RULER(Hsieh et al., 2024), the inability to access and utilize arbitrary information from long contexts remains a critical limitation for many applications, from multi-round chatbots to agent-based systems (Liu et al., 2024; Pan et al., 2025; Guo et al., 2023).
@@ 2 Related Work
Approaches to long-context modeling can be broadly categorized. The first comprises local methods like Sliding Window Attention (Beltagy et al., 2020; Xiao et al., 2024). While effective at maintaining perplexity, they are fundamentally limited by a fixed-size local receptive field, preventing access to information beyond their local window. A second category includes recurrent architectures  (Gu & Dao, 2024; Dao & Gu, 2020; Sun et al., 2023; Qin et al., 2023; Peng et al., 2023; Yang et al., 2024). Recent work shows that zero-initialized recurrent models struggle to generalize to longer sequences; it attributes the failure to models entering unexplored states and introduces State Passing and Truncated Backpropagation Through Time (Sutskever, 2013; Williams & Peng, 1990) to improve length generalization, achieving stable perplexity on out-of-distribution lengths (Buitrago & Gu, 2025). However, on challenging retrieval tasks (Hsieh et al., 2024), performance still drops significantly beyond the training length. This is because the entire context history is compressed into a fixed-size state, sacrificing the ability to recall specific, distant information losslessly. The third category seeks to extend full attention by modifying positional encodings (Press et al., 2022; Peng et al., 2024; Chen et al., 2023; Tan et al., 2025). Their training-free extrapolation gains are modest, typically limited to a small multiple of training length.
In pursuit of more scalable extrapolation, chunk-based sparse attention has emerged as a distinct and highly effective paradigm. Landmark Attention (Mohtashami & Jaggi, 2023) pioneered this by using dedicated landmark tokens to represent chunks for retrieval. While innovative, its length exptrapolation capability is limited to about 64x training length. NSA (Yuan et al., 2025) proposed an efficient hardware-aligned implementation of chunk selection but has very limited length extrapolation (Hu et al., 2025a). State-of-the-art training-free extrapolation has been achieved by Differentiable Retrieval-based Transformers (DRT) (Hu et al., 2025b) and RAMba (Hu et al., 2025a), which can generalize up to 1000x their training length on simple retrieval tasks by introducing a learnable hierarchical chunkwise sparse attention for retrieval. These methods also differ in their retrieval frequency, with DRT operating per-chunk while RAMba, NSA, and Landmark retrieve per-token. Our work builds upon the SWA+Hierarchical Sparse Attention (HSA) architecture with per-token retrieval, providing a systematic analysis to dissect and identify the specific architectural components responsible for the extreme generalization capabilities of this emerging paradigm.
@@ 3.1 Random Context Access: The Key Ability to Utilize Full Context
Under this definition, different architectures exist on a spectrum of capability. The standard Transformer, with its full-attention mechanism, perfectly exemplifies RCA by granting unrestricted access to the entire KV cache. In stark contrast, many efficient architectures trade this property for computational gains by imposing structural constraints. These include: locality constraints, in which models like Sliding Window Attention are confined to a local neighborhood; information bottlenecks, created by State Space Models (SSMs) that compress history into a lossy, fixed-size state; and static access patterns (e.g., Child et al. (2019); Beltagy et al. (2020)), where data-independent sparse attention predetermines information pathways, with more details in Apx. B.1. These architectural bottlenecks fundamentally limit a model’s capacity for true long-range reasoning.
@@ 3.2 Dynamic Chunkwise Sparse Attention
Figure 1: SWA+HSA Architecture. Our SWA+HSA model is primarily composed of Transformers with Sliding Window Attention (SWA), with layers divided into upper layers and lower layers. Lower decoders are composed of SWA followed by an FFN. At the architectural midpoint, a chunking layer splits representations from lower decoders into chunks, and encode them by chunk to build a global memory, each chunk of which contains a landmark vectors ($\mathbf{lmk}$) and encoded chunks ($\mathbf{E}$). Upper decoders each contains a local self-attention layer for local context, followed by an HSA to attend selected global memory and then an FFN.
@@ 3.3 Hierarchical Sparse Attention and Model Design
As shown in Figure 1, our SWA+HSA model treats the global memory in a chunkwise approach. During decoding, the chunking layer gathers hidden representations into a buffer and encodes them as complete chunks. To maintain causality, the chunk containing the current token, as well as all subsequent chunks, are not accessible for retrieval. Following the notation of Sun et al. (2023), we use indices with $\cdot$] to indicate that it is indexed by chunk rather than by token. For instance, $\mathbf{H}_{[t]}:=\mathbf{h}_{t\cdot C+1:(t+1)\cdot C}\in\mathbb{R}^{C\times d}$ is the chunked hidden states. At chunk selection layer, a query vector $\mathbf{q_{t}}\in\mathbb{R}^{d_{\text{Rtrv}}}$ computes a dot-product similarity score on landmark $\mathbf{lmk}_{[i]}\in\mathbb{R}^{d_{\text{Rtrv}}}$, formulated as $s_{t,i}=\mathbf{q_{t}}\cdot\mathbf{{lmk}}_{[i]}\in\mathbb{R}$. The indices of the top-$N$ chunks are then selected, forming set $\mathcal{I}_{t}$.
A critical design choice lies in how these selected chunks are weighted, as this determines how the retrieval signal is backpropagated. The weighting function $w_{t,i}$ is zero for all non-selected chunks. For the selected chunks, the weights are assigned according to distinct strategies for NSA (Eq. 1a), GCA (Eq. 1b), and HSA (Eq. 1c). Let $\mathcal{S}_{t}=\{s_{t,i}\}_{i\in\mathcal{I}_{t}}$.
This comparison highlights a clear progression: from the implicit binary weighting of NSA, to the proportional softmax weighting in GCA (in Hu et al. (2025b)), and finally to the recency-biased allocation via stick-breaking process  [^1] Specifically, let $\mathcal{I}^{\prime}_{t}=\text{sort}(\mathcal{I}_{t})$ be the selected chunk indices sorted by their scores, then $w_{t,k}=\sigma(s_{t,\mathcal{I}^{\prime}_{t,k}})\prod_{j<k}(1-\sigma(s_{t,\mathcal{I}^{\prime}_{t,j}}))$ where $\sigma$ is the sigmoid function. (Tan et al., 2025) used in HSA (our work and Hu et al. (2025a)).
@@ 4.1 Chunk Information preprocess
Chunked Hidden States: Disentangling Representations for Retrieval and Prediction. The hidden state $\mathbf{h}^{\frac{L}{2}}_{t}$ must be immediately useful for autoregressive next token prediction, while also serving as a durable memory for future retrieval. Considering the dual purposes of $\mathbf{h}^{\frac{L}{2}}_{t}$, we posit that simply operating on these raw hidden states conflates these two distinct functional objectives. The introduction of non-linear layers aims to disentangle these two roles. By processing the raw states, these components can produce refined representations specifically optimized for retrieval, separate from the immediate demands of next-token prediction. A RMSNorm provides a simple mechanism to introduce non-linearity, while the Encoder takes this a step further by explicitly learning to re-encode the chunk’s information.
@@ 5 Experiments
Figure 3: BabiLong Evaluation Results: The best SWA+HSA model (ours) demonstrates robust long-context extrapolation, maintaining high accuracy up to 8M tokens. In stark contrast, baseline models exhibit clear limitations. Full-attention methods collapse to near-zero accuracy shortly after their training length. Prior sparse methods like Landmark Attention show better generalization but also fail decisively at longer contexts ($\geq$64k). Sliding Window Attention plateaus at a low accuracy consistent with random guessing, while Mamba2’s performance degrades substantially.
Table 2: RULER evaluation results. Our best SWA+HSA model maintains high accuracy up to 32M tokens, significantly outperforming all baselines in long-context scenarios. In Model Specs, ‘Enc’ denotes the number of encoder layers, ‘CLS’ indicates the use of a CLS token, and ‘Bypass’ refers to the Bypassing Residual Path.
@@
SWA+HSA 0 no no 72.28 21.78 3.96 32.67 61.39 6.93 1.98 23.43 61.80 1.12 0.00 20.97
SWA+HSA 1 no no 72.73 58.18 72.73 67.88 69.09 34.55 58.18 53.94 70.91 47.27 43.64 53.94
SWA+HSA 1 yes no 63.64 40.00 54.55 52.73 69.09 38.18 49.09 52.12 54.55 21.82 30.91 35.76
SWA+HSA 2 no no 63.37 63.37 72.28 66.34 58.42 48.51 66.34 57.76 67.42 42.70 70.79 60.30
SWA+HSA 2 yes no 83.17 46.53 82.18 70.63 78.22 28.71 80.20 62.38 82.02 24.72 61.80 56.18
SWA+HSA 0 no yes 80.20 64.36 45.54 63.37 71.29 46.53 15.84 44.55 82.02 31.46 3.37 38.95
SWA+HSA 1 no yes 69.09 56.36 45.45 56.97 58.18 30.91 10.91 33.33 63.64 36.36 12.73 37.58
SWA+HSA 1 yes yes 89.09 69.09 80.00 79.39 72.73 78.18 76.36 75.76 85.45 69.09 69.09 74.54
SWA+HSA 2 no yes 85.15 74.26 74.26 77.89 78.22 65.35 65.35 69.64 82.02 66.29 41.57 63.29
SWA+HSA 2 yes yes 88.12 76.24 93.07 85.81 79.21 77.23 90.10 82.18 82.02 68.54 88.76 79.77
We organize our experiments as follows. Section 5.1 describes the experimental setup, including model architecture, training details, benchmarks, and baselines. Section 5.2 presents length extrapolation results on BabiLong and RULER, demonstrating the superiority of our SWA+HSA architecture. Section 5.3 provides an in-depth analysis of the retrieval mechanism, disentangling retrieval accuracy from information integration. Section 5.4 investigates the role of sparsity in enabling extreme-length generalization. Finally, Section 5.5 validates the scalability and general capabilities of our approach.
5.1 Experimental Setup
Our experiments are conducted on a series of models based on the SWA+HSA architecture detailed in Figure 1. All model variants contain approximately 240M parameters and are trained with a context length of 4K. Unlike the original DRT (Hu et al., 2025b) which uses Grouped Cross Attention (GCA) for per-chunk retrieval, our architecture employs HSA with a stick-breaking weighting strategy (Eq. 1c) to enable per-token retrieval, providing greater flexibility for dynamic access to long-range information.
Key hyperparameters for this setup are fixed across experiments. The sliding window size for the lower layers is 512 tokens, and the HSA chunk size is 64. Unless otherwise noted, we use a Top-K of 8 for chunk selection, giving the HSA module an effective receptive field of 512 tokens. A comprehensive description of our training recipes and further hyperparameter details is provided in Appendix C.
To complement this and more directly assess a model’s core retrieval ability in isolation, we also use RULER. From RULER, we select three representative sub-tasks: Single-Needle (S-N); Multi-Query Needle (MQ-N), where 2 of 6 key-value pairs are queried; and Variable Tracking (VT), which tests transitive assignment chains. Collectively, these tasks provide a targeted evaluation of a model’s capacity for precise, long-range information retrieval and tracking simple logical chains. As RULER requires accurately retrieving several tokens, it provides a strict test of models’ random context access ability.
Baselines. We compare our method against several baselines. Our primary full-attention baseline is Llama with YaRN and Mamba2 with State Passing (Buitrago & Gu, 2025), which is trained on a 4k context length. To ensure a fair comparison with other efficient attention mechanisms, we include Landmark Attention, configured to match our model’s training FLOPs during training by using a sequence length of 756. Furthermore, to benchmark against methods with a similar computational budget, we include Stick-breaking Attention and Sliding Window Attention. Both of these are trained with a 1k context window to keep the receptive field same to our approach.
5.2 Length Extrapolation Results
@@
As shown in Figure 3, the results on BabiLong highlight the stark differences in extrapolation capability across architectures. The benchmark’s classification-like nature establishes a random-guess baseline around 0.3 accuracy, a level to which Sliding Window Attention predictably converges due to its limited context. Other baselines also fail to generalize, with full-attention and Landmark Attention collapsing at various points beyond their training regimes. Crucially, the figure also reveals the importance of our key design choices within the SWA+HSA architecture. The model variants with a learnable chunk encoder consistently outperform those without.
@@
To further validate these findings and provide a more granular analysis of retrieval performance, we turn to the RULER benchmark, with detailed results in Table 2. The trends observed here strongly corroborate our conclusions from BabiLong. First, comparing our SWA+HSA models against the baselines reveals a significant performance gap in extrapolation. Llama with Yarn (Llama-Yarn), despite its strong in-domain performance at 4K, fails catastrophically at 32K, with their scores dropping to zero. Stick-breaking Attention (Stick-Break) and Landmark Attention demonstrate some extrapolation ability but exhibit significant performance decay and are consistently outperformed by our models on longer contexts.
@@
Second, while all our model variants showcase robust long-context performance, the results also highlight a clear performance hierarchy among our different architectural configurations. Our full model (Enc=2, CLS=yes, Bypass=yes) consistently achieves state-of-the-art results. It not only secures the highest average score of 87.92% at the 4K training length but also degrades remarkably gracefully, maintaining an average score of 79.77% even at an extreme context length of 32M. This indicates that the specific combination of our proposed components is critical for unlocking maximum performance, an interaction we will analyze in detail in the subsequent section.